# Modeling the dynamics of an oil drop driven by a surface acoustic wave in the underlying substrate

### v3

@@ 5 Conclusions
In this work, we consider the problem of a spreading silicone oil drop driven by a surface acoustic wave (SAW) propagating in the supporting substrate. Our particular focus is on a theoretical framework that describes drops of millimetric thickness, for which the main driving mechanism is the Eckart streaming. Therefore, we focus a different regime from that relevant for the much thinner drops considered by Rezk et al. (2012, 2014), for which Schlichting and Rayleigh streaming effects are dominant. To the best of our knowledge, modeling the dynamics of millimetric drops under the action of SAW and capillary stresses has not been considered so far in the literature. We also present selected experimental results for the qualitative validation of our model.
@@
While not all details of the experimental results are fully understood, in general, we find reasonable agreement between the experimental and theoretical results, supporting further the basic premises underlying the developed acousto-fluidic model. In particular, our model provides good qualitative predictions for drop shapes, spreading speeds, and the dependence of the results on the SAW displacement amplitude at the solid surface (the SAW intensity). Both theoretical and experimental results show asymmetric drop shapes with a thin trailing film left behind, and both appear to reach a constant spreading speed asymptotically in time. The latter finding motivated us to formulate a simplified traveling wave model. The resulting analytical solution predicts an approximately linear increase in the drop thickness and a quadratic increase of the spreading speed with the acoustic intensity (SAW normal displacement amplitude at the solid surface), $A$. Such predicted trends are consistent with those of the full theoretical model and the experiments.
While many features of the experimental results have been rationalized using our theoretical models and simulations, much remains to be done. Further investigation is required to shed light on the features of dynamic drop profiles under SAW excitation. The transition from three-dimensional (3D) to quasi-two-dimensional (2D) dynamics is also intriguing: under strong SAW excitation, a drop spreads along the path of the SAW and may be described to leading order using the 2D model given here. However, at weak SAW excitation, where the drop also spreads radially due to the relaxation of capillary stress at the free drop surface, the drop dynamics is inherently of the 3D type. Hence, 3D aspects of drop dynamics, in particular in cases where the contributions to flow from capillary and SAW stresses are comparable, remain to be simulated and understood.
@@ Appendix A First-Order Velocity Components and SAW Force
As mentioned in the main text, although we deal with imaginary equations, it is understood that the real part of these complex equations is taken in results for real quantities. This includes the first-order velocity field, $\mathbf{v}_{1}$, given by Eq. (3.1), and the acoustic force on the induced steady streaming flow, $\mathbf{F}_{s}$, given by Eq. (17), which we now discuss in detail.
@@
$V_{x}$ $=$ $\frac{1}{\Delta}\left[k_{\rm s,i}\cos\left(\frac{\theta}{2}+k_{\rm s,r}x+z\Delta\sin\left(\frac{\theta}{2}\right)-\omega t\right)+k_{\rm s,r}\sin\left(\frac{\theta}{2}+k_{\rm s,r}x+z\Delta\sin\left(\frac{\theta}{2}\right)-\omega t\right)\right],$ (42)
$V_{z}$ $=$ $\cos\left(k_{\rm s,r}x+z\Delta\sin\left(\frac{\theta}{2}\right)-\omega t\right),$ (43)
@@
$\Delta$ $=\sqrt{\left|\kappa_{\rm s}^{2}-\kappa^{2}\right|}=\sqrt[4]{\left(k_{\rm s,i}^{2}-k_{\rm s,r}^{2}+k_{\rm r}^{2}-k_{\rm i}^{2}\right)^{2}+(2k_{\rm s,i}k_{\rm s,r}-2k_{\rm i}k_{\rm r})^{2}},$ (44)
@@
Turning next to $\mathbf{F}_{s}$, in order to find the real part we substitute Eq. (41) with the subsequently defined parameters into Eq. (17), and find explicit expressions for the acoustic force components in the form,
$F_{s,x}$ $=$ $C_{x}P_{0}\,e^{-2(k_{\rm s,i}x+K_{z}z)},$ (47)
$F_{s,z}$ $=$ $C_{z}P_{0}\,e^{-2(k_{\rm s,i}x+K_{z}z)},$ (48)
@@
$C_{x}$ $=$ $\frac{k_{\rm s,i}k_{\rm s,r}^{2}+\Delta\cos\left(\frac{\theta}{2}\right)k_{\rm s,i}K_{z}+k_{\rm s,i}^{3}+\Delta\sin\left(\frac{\theta}{2}\right)k_{\rm s,r}K_{z}}{\Delta^{2}},$ (49)
$C_{z}$ $=$ $\frac{\cos\left(\frac{\theta}{2}\right)k_{\rm s,i}^{2}+\sin\left(\frac{\theta}{2}\right)k_{\rm s,i}k_{\rm s,r}+\Delta\,K_{z}}{\Delta}.$ (50)
We point out that $\mathbf{F}_{s}$ is a non–conservative force; it cannot be written as the gradient of a potential, except for a special choice of parameter values that satisfy $k_{\rm s,i}C_{z}=K_{z}C_{x}$ (see Eqs. (47) and (48)).
@@ B.1 Oil (PDMS)
In this section, we take a closer look at the acoustic force $\mathbf{F}_{s}$, given by Eq. (17), on the induced steady streaming flow. We first note that, for our parameter values, we can define the small quantities
$\varepsilon_{s}=\frac{k_{\rm s,i}}{k_{\rm s,r}}\qquad\text{and}\qquad\varepsilon=\frac{k_{\rm i}}{k_{\rm r}}\approx\frac{\beta}{2}.$ (51)
These two parameters $\varepsilon_{s}$ and $\varepsilon$ are always small, and of the same order of magnitude, for the PDMS used here (see Eq. (12) and the surrounding discussion). To approximate the SAW force, we can expand the auxiliary variables and SAW force coefficients, given in Eqs. (44), (45), (46), (49) and (50), in terms of $\varepsilon_{s}$ and $\varepsilon$. Retaining only linear terms in the expansions (neglecting any terms quadratic in $\varepsilon_{s},\varepsilon$ that are of higher order than those retained) yields the following expressions
$\Delta$$\approx$ $k_{\rm s,r}\alpha_{1},$
$\theta$$\approx$ $\pi-\frac{2}{\alpha_{1}^{2}}(1+\alpha_{1}^{2})\varepsilon+\frac{2}{\alpha_{1}^{2}}\varepsilon_{s},$
$K_{z}$$\approx$ $\frac{k_{\rm s,r}}{\alpha_{1}}(1+\alpha_{1}^{2})\varepsilon-\frac{k_{\rm s,r}}{\alpha_{1}}\varepsilon_{s},$ (52)
$C_{x}$$\approx$ $\frac{k_{\rm s,r}(1+\alpha_{1}^{2})}{\alpha_{1}^{2}}\varepsilon,$
@@
$\alpha_{1}=\sqrt{\frac{k^{2}}{k_{\rm s,r}^{2}}-1}=\sqrt{\frac{c_{\rm s}^{2}}{c^{2}}-1},$ (53)
since $k_{\rm s,r}=\omega/c_{\rm s}$. This coefficient yields the Rayleigh angle $\theta_{R}$ as: $\tan\theta_{R}=1/\alpha_{1}$.
@@
$F_{s,x}$ $\approx\frac{\rho_{0}A^{2}\omega^{2}(1+\alpha_{1})^{2}k_{\rm s,r}}{\alpha_{1}^{2}}\varepsilon\ e^{-2(k_{\rm s,i}x+K_{z}z)},$ (54)
$F_{s,z}$ $\approx\alpha_{1}F_{s,x}.$ (55)
The approximations made in Eqs. (52) lead to an attenuation factor $K_{z}$ that is approximately $12$% smaller and force coefficients $C_{x}$ and $C_{z}$ that are approximately $2$% bigger and $6$% smaller, respectively, with respect to the quantities in the exact expressions given in Eqs. (46),  (49), and (50). Note that the approximate force components obtained in Eqs. (54), (55) (also) do not yield a conservative force, since $K_{z}\neq\alpha_{1}k_{\rm s,i}$.
@@ B.2 Water
Next, we consider modeling a lower viscosity fluid such as water, characterized by a similar density and phase velocity to oil (PDMS) so that the imaginary component of the wavenumber in the solid is similar (see Eq. (11)). However, the large decrease in viscosity in the case of water leads to a much reduced value of the imaginary wavenumber in the liquid, $k_{\rm i}$ (see Eq. (7)). Thus for water, our assumption that $\varepsilon$ and $\varepsilon_{s}$ are of the same order of magnitude fails; in fact, here we have $\varepsilon_{s}^{2}\sim\varepsilon$ so that we need to also retain terms that are quadratic in $\varepsilon_{s}$; the next largest terms in the expansion are those proportional to $\varepsilon_{s}^{3}$ which we also include for the purpose of comparison to the approximation in the case of neglecting viscous dissipation which is discussed later in Sec. B.3. Note that we neglect any cross terms that may appear in the expansions as they are higher order than the terms retained.
@@
$\Delta$ $\approx k_{\rm s,r}\alpha_{1}+\varepsilon_{s}^{2}\bigg{(}\frac{k_{\rm s,r}}{\alpha_{1}^{3}}+\frac{k_{\rm s,r}}{2\alpha_{1}}\bigg{)},$
$\theta$ $\approx-\pi+\frac{2}{\alpha_{1}^{2}}\varepsilon_{s}-\frac{2(1+\alpha_{1}^{2})}{\alpha_{1}^{2}}\varepsilon,$
$K_{z}$ $\approx\frac{k_{\rm s,r}}{\alpha_{1}}\varepsilon_{s}-\frac{k_{\rm s,r}(1+\alpha_{1}^{2})}{\alpha_{1}}\varepsilon,$ (56)
$C_{x}$ $\approx\frac{k_{\rm s,r}(1+\alpha_{1}^{2})}{\alpha_{1}^{2}}\bigg{(}\varepsilon+\frac{\varepsilon_{s}^{3}}{\alpha_{1}^{2}}\bigg{)},$
$C_{z}$ $\approx\frac{k_{\rm s,r}(1+\alpha_{1}^{2})}{\alpha_{1}}\bigg{(}-\varepsilon+\frac{\varepsilon_{s}^{3}}{\alpha_{1}^{4}}\bigg{)}.$
We pause to discuss the difference between the approximated complex arguments in the case of oil and water. Note that the two expressions for $\theta$ (Eqs. (52) and (56)) have the same first-order correction terms with respect to the small parameters $\varepsilon$ and $\varepsilon_{s}$; however, the sign of the sum of these two terms is opposite for the case of PDMS (negative) and water (positive) - which is due to the much reduced value of the imaginary wavenumber in the liquid, $k_{\rm i}$, in the case of water as compared to PDMS (see Eqs. (6) and (8)). The result of this sign change has the effect of moving the complex argument from quadrant II (in the case of PDMS) to quadrant III (in the case of water); as a result, the approximation for the sine function in the SAW force coefficients changes signs and thus the overall value of the SAW force coefficients is much reduced (see Eqs. (49) and (50)).
@@
$F_{s,x}$ $\approx\frac{\rho_{0}A^{2}\omega^{2}(1+\alpha_{1})^{2}k_{\rm s,r}}{\alpha_{1}^{2}}\varepsilon\ e^{-2(k_{\rm s,i}x+K_{z}z)},$ (57)
$F_{s,z}$ $\approx-\alpha_{1}F_{s,x}.$ (58)
@@ B.3 Neglecting Viscous Dissipation
$\Delta$$\approx$ $k_{\rm s,r}\alpha_{1},$
$\theta$$\approx$ $-\pi+\frac{2}{\alpha_{1}^{2}}\varepsilon_{s},$
$K_{z}$$\approx$ $\frac{k_{\rm s,r}}{\alpha_{1}}\varepsilon_{s},$ (59)
$C_{x}$$\approx$ $\frac{k_{\rm s,r}(1+\alpha_{1}^{2})}{\alpha_{1}^{4}}\varepsilon_{s}^{3},$
@@
$F_{s,x}$ $\approx\frac{\rho A^{2}\omega^{2}(1+\alpha_{1}^{2})}{\alpha_{1}^{4}}k_{\rm s,r}\varepsilon_{s}^{3}e^{-2k_{\rm s,i}(x+\frac{1}{\alpha_{1}}z)},$ (60)
$F_{s,z}$ $\approx\frac{1}{\alpha_{1}}F_{s,x},$ (61)
which can be rewritten in terms of a potential $\phi_{s}$, so that $\vec{F}_{s}=-\nabla\phi_{s}$ is a conservative volume force with
$\phi_{s}=\frac{\rho A^{2}\omega^{2}(1+\alpha_{1}^{2})}{2\alpha_{1}^{4}}\varepsilon_{s}^{2}e^{-2k_{\rm s,i}(x+\frac{1}{\alpha_{1}}z)}~{}.$ (62)
We note that without viscous dissipation the force components are five orders of magnitude smaller than when the viscosity is considered, compare for example Eqs. (49), (50) to Eq. (59) with the given parameter values. As a result, the force in Eqs. (60) and (61) has no effects on drop dynamics on the timescales investigated. Such finding inspires comparison with the results presented in Shiokawa et al. (1990, 1989), and given in terms of our variables as
$\phi_{\rm shio}=\frac{1+\alpha_{1}^{2}}{2}\rho A^{2}\omega^{2}e^{-2k_{\rm s,i}(x+\alpha_{1}z)}.$ (63)
Note that, for the considered parameter values, this expression leads to a force that is approximately three orders of magnitude larger compared to what is found from our model, i.e., Eq. (62). Further, our model predicts a different attenuation factor in $z$ as well as a different force prefactor than what is found in Shiokawa et al. (1990, 1989). In those works, the authors find that $K_{z}=\alpha_{1}k_{\rm s,i}$, as opposed to our model which yields $K_{z}\approx{k_{\rm s,r}\varepsilon_{s}/\alpha_{1}}=k_{\rm s,i}/\alpha_{1}$. From the physical point of view, one expects larger viscous dissipation at a given $z$ (distance from the substrate, into the liquid), when the Rayleigh angle, $\theta_{R}$, increases and tends to $\pi/2$ (see Fig. 4, where the angle is measured clockwise from the normal to the substrate). This is because the leaky wave travels a longer distance along the ray (for given $z$) for larger $\theta_{R}$, and thus its amplitude is more attenuated. For example, if $c\rightarrow c_{s}$, we have $\theta_{R}\rightarrow\pi/2$ and $\alpha_{1}=1/\tan\theta_{R}\rightarrow 0$ (from Snell’s law, we have $\sin(\theta_{R})=c/c_{s}<1$). Since $K_{z}$ must increase for $\theta_{R}\rightarrow\pi/2$, we have $\tan\theta_{R}\rightarrow\infty$, and then one expects $K_{z}\propto 1/\alpha_{1}\rightarrow\infty$, and not $K_{z}\propto\alpha_{1}\rightarrow 0$ as in Shiokawa et al. (1990, 1989).

### v4

@@ 5 Conclusions
In this work, we consider the problem of a spreading silicone oil drop driven by a surface acoustic wave (SAW) propagating in the supporting substrate. Our particular focus is on a theoretical framework that describes drops of millimetric thickness, for which the main driving mechanism is the Eckart streaming. Therefore, we focus on a different regime from that relevant for the much thinner drops considered by Rezk et al. (2012, 2014), for which Schlichting and Rayleigh streaming effects are dominant. To the best of our knowledge, modeling the dynamics of millimetric drops under the action of SAW and capillary stresses has not been considered so far in the literature. We also present selected experimental results for the qualitative validation of our model.
@@
While not all details of the experimental results are fully understood, in general, we find reasonable agreement between the experimental and theoretical results, further supporting the basic premises underlying the developed acousto-fluidic model. In particular, our model provides good qualitative predictions for drop shapes, spreading speeds, and the dependence of the results on the SAW displacement amplitude at the solid surface (the SAW intensity). Both theoretical and experimental results show asymmetric drop shapes with a thin trailing film left behind, and both appear to reach a constant spreading speed asymptotically in time. The latter finding motivated us to formulate a simplified traveling wave model. The resulting analytical solution predicts an approximately linear increase in the drop thickness and a quadratic increase of the spreading speed with the acoustic intensity (SAW normal displacement amplitude at the solid surface), $A$. Such predicted trends are consistent with those of the full theoretical model and the experiments.
While many features of the experimental results have been rationalized using our theoretical models and simulations, much remains to be done. Further investigation is required to shed light on the features of dynamic drop profiles under SAW excitation. The transition from three-dimensional (3D) to quasi-two-dimensional (2D) dynamics is also intriguing: under strong SAW excitation, a drop spreads primarily along the path of the SAW and may be described to leading order using the 2D model given here. However, at weak SAW excitation, where the drop also spreads radially due to the relaxation of capillary stress at the free drop surface, the drop dynamics is inherently of the 3D type. Hence, 3D aspects of drop dynamics, in particular in cases where the contributions to flow from capillary and SAW stresses are comparable, remain to be simulated and understood.
@@ Appendix A First-Order Velocity Components and SAW Force
As mentioned in the main text, although we deal with imaginary equations in the model derivation, it is understood that the real part of these complex equations is taken as the result for physically-relevant real quantities. This includes the first-order velocity field, $\mathbf{v}_{1}$, given by Eq. (3.1), and the acoustic force on the induced steady streaming flow, $\mathbf{F}_{\rm s}$, given by Eq. (17), which we now discuss in detail.
@@
$V_{x}$ $=$ $\frac{1}{\Delta}\left[k_{\rm s,{\rm i}}\cos\left(\frac{\theta}{2}+k_{\rm s,{\rm r}}x+z\Delta\sin\left(\frac{\theta}{2}\right)-\omega t\right)+k_{\rm s,{\rm r}}\sin\left(\frac{\theta}{2}+k_{\rm s,{\rm r}}x+z\Delta\sin\left(\frac{\theta}{2}\right)-\omega t\right)\right],$ (43)
$V_{z}$ $=$ $\cos\left(k_{\rm s,{\rm r}}x+z\Delta\sin\left(\frac{\theta}{2}\right)-\omega t\right),$ (44)
@@
$\Delta$ $=\sqrt{\left|\kappa_{\rm s}^{2}-\kappa^{2}\right|}=\sqrt[4]{\left(k_{\rm s,{\rm i}}^{2}-k_{\rm s,{\rm r}}^{2}+k_{\rm r}^{2}-k_{\rm i}^{2}\right)^{2}+(2k_{\rm s,{\rm i}}k_{\rm s,{\rm r}}-2k_{\rm i}k_{\rm r})^{2}},$ (45)
@@
Turning next to $\mathbf{F}_{\rm s}$, in order to find the real part we substitute Eq. (42) with the subsequently defined parameters into Eq. (17), and find explicit expressions for the acoustic force components in the form,
$F_{{\rm s},x}$ $=$ $C_{x}P_{0}\,e^{-2(k_{\rm s,{\rm i}}x+K_{z}z)},$ (48)
$F_{{\rm s},z}$ $=$ $C_{z}P_{0}\,e^{-2(k_{\rm s,{\rm i}}x+K_{z}z)},$ (49)
@@
$C_{x}$ $=$ $\frac{k_{\rm s,{\rm i}}k_{\rm s,{\rm r}}^{2}+\Delta\cos\left(\frac{\theta}{2}\right)k_{\rm s,{\rm i}}K_{z}+k_{\rm s,{\rm i}}^{3}+\Delta\sin\left(\frac{\theta}{2}\right)k_{\rm s,{\rm r}}K_{z}}{\Delta^{2}},$ (50)
$C_{z}$ $=$ $\frac{\cos\left(\frac{\theta}{2}\right)k_{\rm s,{\rm i}}^{2}+\sin\left(\frac{\theta}{2}\right)k_{\rm s,{\rm i}}k_{\rm s,{\rm r}}+\Delta\,K_{z}}{\Delta}.$ (51)
We point out that $\mathbf{F}_{{\rm s}}$ is a non–conservative force; it cannot be written as the gradient of a potential, except for a special choice of parameter values that satisfy $k_{\rm s,{\rm i}}C_{z}=K_{z}C_{x}$ (see Eqs. (48) and (49)).
@@ B.1 Oil (PDMS)
In this section, we take a closer look at the acoustic force $\mathbf{F}_{\rm s}$, given by Eq. (17), on the induced steady streaming flow. We first note that, for our parameter values, we can define the small quantities
$\varepsilon_{\rm s}=\frac{k_{\rm s,{\rm i}}}{k_{\rm s,{\rm r}}}\qquad\text{and}\qquad\varepsilon=\frac{k_{\rm i}}{k_{\rm r}}\approx\frac{\beta}{2}.$ (52)
These two parameters $\varepsilon_{\rm s}$ and $\varepsilon$ are always small, and of the same order of magnitude, for the PDMS used here (see Eq. (12) and the surrounding discussion). To approximate the SAW force, we can expand the auxiliary variables and SAW force coefficients, given in Eqs. (45), (46), (47), (50) and (51), in terms of $\varepsilon_{\rm s}$ and $\varepsilon$. Retaining only linear terms in the expansions (neglecting any terms quadratic in $\varepsilon_{\rm s},\varepsilon$ that are of higher order than those retained) yields the following expressions
$\Delta$ $\approx$ $k_{\rm s,{\rm r}}\alpha_{1},$
$\theta$ $\approx$ $\pi-\frac{2}{\alpha_{1}^{2}}(1+\alpha_{1}^{2})\varepsilon+\frac{2}{\alpha_{1}^{2}}\varepsilon_{\rm s},$
$K_{z}$ $\approx$ $\frac{k_{\rm s,{\rm r}}}{\alpha_{1}}(1+\alpha_{1}^{2})\varepsilon-\frac{k_{\rm s,{\rm r}}}{\alpha_{1}}\varepsilon_{\rm s},$ (53)
$C_{x}$ $\approx$ $\frac{k_{\rm s,{\rm r}}(1+\alpha_{1}^{2})}{\alpha_{1}^{2}}\varepsilon,$
@@
$\alpha_{1}=\sqrt{\frac{k^{2}}{k_{\rm s,{\rm r}}^{2}}-1}=\sqrt{\frac{c_{\rm s}^{2}}{c^{2}}-1},$ (54)
since $k_{\rm s,{\rm r}}=\omega/c_{\rm s}$. This coefficient yields the Rayleigh angle $\theta_{R}$ as: $\tan\theta_{R}=1/\alpha_{1}$.
@@
$F_{{\rm s},x}$ $\approx\frac{\rho_{0}A^{2}\omega^{2}(1+\alpha_{1})^{2}k_{\rm s,{\rm r}}}{\alpha_{1}^{2}}\varepsilon\ e^{-2(k_{\rm s,{\rm i}}x+K_{z}z)},$ (55)
$F_{{\rm s},z}$ $\approx\alpha_{1}F_{{\rm s},x}.$ (56)
The approximations made in Eqs. (53) lead to an attenuation factor $K_{z}$ that is approximately 1212% smaller and force coefficients $C_{x}$ and $C_{z}$ that are approximately 22% bigger and 66% smaller, respectively, with respect to the quantities in the exact expressions given in Eqs. (47),  (50), and (51). Note that the approximate force components obtained in Eqs. (55), (56) (also) do not yield a conservative force, since $K_{z}\neq\alpha_{1}k_{\rm s,{\rm i}}$.
@@ B.2 Water
Next, we consider modeling a lower viscosity fluid such as water, characterized by a similar density and phase velocity to oil (PDMS) so that the imaginary component of the wavenumber in the solid is similar (see Eq. (11)). However, the large decrease in viscosity in the case of water leads to a much reduced value of the imaginary wavenumber in the liquid, $k_{\rm i}$ (see Eq. (7)). Thus for water, our assumption that $\varepsilon$ and $\varepsilon_{\rm s}$ are of the same order of magnitude fails; in fact, here we have $\varepsilon_{\rm s}^{2}\sim\varepsilon$ so that we need to also retain terms that are quadratic in $\varepsilon_{\rm s}$; the next largest terms in the expansion are those proportional to $\varepsilon_{\rm s}^{3}$ which we also include for the purpose of comparison to the approximation in the case of neglecting viscous dissipation which is discussed later in Sec. B.3. Note that we neglect any cross terms that may appear in the expansions as they are higher order than the terms retained.
@@
$\Delta$ $\approx k_{\rm s,{\rm r}}\alpha_{1}+\varepsilon_{\rm s}^{2}\bigg(\frac{k_{\rm s,{\rm r}}}{\alpha_{1}^{3}}+\frac{k_{\rm s,{\rm r}}}{2\alpha_{1}}\bigg),$
$\theta$ $\approx-\pi+\frac{2}{\alpha_{1}^{2}}\varepsilon_{\rm s}-\frac{2(1+\alpha_{1}^{2})}{\alpha_{1}^{2}}\varepsilon,$
$K_{z}$ $\approx\frac{k_{\rm s,{\rm r}}}{\alpha_{1}}\varepsilon_{\rm s}-\frac{k_{\rm s,{\rm r}}(1+\alpha_{1}^{2})}{\alpha_{1}}\varepsilon,$ (57)
$C_{x}$ $\approx\frac{k_{\rm s,{\rm r}}(1+\alpha_{1}^{2})}{\alpha_{1}^{2}}\bigg(\varepsilon+\frac{\varepsilon_{\rm s}^{3}}{\alpha_{1}^{2}}\bigg),$
$C_{z}$ $\approx\frac{k_{\rm s,{\rm r}}(1+\alpha_{1}^{2})}{\alpha_{1}}\bigg(-\varepsilon+\frac{\varepsilon_{\rm s}^{3}}{\alpha_{1}^{4}}\bigg).$
We pause to discuss the difference between the approximated complex arguments in the case of oil and water. Note that the two expressions for $\theta$ (Eqs. (53) and (57)) have the same first-order correction terms with respect to the small parameters $\varepsilon$ and $\varepsilon_{\rm s}$; however, the sign of the sum of these two terms is opposite for the case of PDMS (negative) and water (positive) - which is due to the much reduced value of the imaginary wavenumber in the liquid, $k_{\rm i}$, in the case of water as compared to PDMS (see Eqs. (6) and (8)). The result of this sign change has the effect of moving the complex argument from quadrant II (in the case of PDMS) to quadrant III (in the case of water); as a result, the approximation for the sine function in the SAW force coefficients changes sign and thus the overall value of the SAW force coefficients is much reduced (see Eqs. (50) and (51)).
@@
$F_{{\rm s},x}$ $\approx\frac{\rho_{0}A^{2}\omega^{2}(1+\alpha_{1})^{2}k_{\rm s,{\rm r}}}{\alpha_{1}^{2}}\varepsilon\ e^{-2(k_{\rm s,{\rm i}}x+K_{z}z)},$ (58)
$F_{{\rm s},z}$ $\approx-\alpha_{1}F_{{\rm s},x}.$ (59)
@@ B.3 Neglecting Viscous Dissipation
$\Delta$ $\approx$ $k_{\rm s,{\rm r}}\alpha_{1},$
$\theta$ $\approx$ $-\pi+\frac{2}{\alpha_{1}^{2}}\varepsilon_{\rm s},$
$K_{z}$ $\approx$ $\frac{k_{\rm s,{\rm r}}}{\alpha_{1}}\varepsilon_{\rm s},$ (60)
$C_{x}$ $\approx$ $\frac{k_{\rm s,{\rm r}}(1+\alpha_{1}^{2})}{\alpha_{1}^{4}}\varepsilon_{\rm s}^{3},$
@@
$F_{{\rm s},x}$ $\approx\frac{\rho A^{2}\omega^{2}(1+\alpha_{1}^{2})}{\alpha_{1}^{4}}k_{\rm s,{\rm r}}\varepsilon_{\rm s}^{3}e^{-2k_{\rm s,{\rm i}}(x+\frac{1}{\alpha_{1}}z)},$ (61)
$F_{{\rm s},z}$ $\approx\frac{1}{\alpha_{1}}F_{{\rm s},x},$ (62)
which can be rewritten in terms of a potential $\phi_{\rm s}$, so that $\vec{F}_{\rm s}=-\nabla\phi_{\rm s}$ is a conservative volume force with
$\phi_{\rm s}=\frac{\rho A^{2}\omega^{2}(1+\alpha_{1}^{2})}{2\alpha_{1}^{4}}\varepsilon_{\rm s}^{2}e^{-2k_{\rm s,{\rm i}}(x+\frac{1}{\alpha_{1}}z)}~.$ (63)
We note that without viscous dissipation, the force components are five orders of magnitude smaller than when the viscosity is considered, compare for example Eqs. (50), (51) to Eq. (60) with the given parameter values. As a result, the force in Eqs. (61) and (62) has no effect on drop dynamics on the timescales investigated. Such a finding inspires comparison with the results presented in Shiokawa et al. (1990, 1989), and given in terms of our variables as
$\phi_{\rm shio}=\frac{1+\alpha_{1}^{2}}{2}\rho A^{2}\omega^{2}e^{-2k_{\rm s,{\rm i}}(x+\alpha_{1}z)}.$ (64)
Note that, for the considered parameter values, this expression leads to a force that is approximately three orders of magnitude larger compared to what is found from our model, i.e., Eq. (63). Further, our model predicts a different attenuation factor in $z$ as well as a different force prefactor than what is found in Shiokawa et al. (1990, 1989). In those works, the authors find that $K_{z}=\alpha_{1}k_{\rm s,{\rm i}}$, as opposed to our model which yields $K_{z}\approx{k_{\rm s,{\rm r}}\varepsilon_{\rm s}/\alpha_{1}}=k_{\rm s,{\rm i}}/\alpha_{1}$. From the physical point of view, one expects larger viscous dissipation at a given $z$ (distance from the substrate, into the liquid), when the Rayleigh angle, $\theta_{R}$, increases and tends to $\pi/2$ (see Fig. 5, where the angle is measured clockwise from the normal to the substrate). This is because the leaky wave travels a longer distance along the ray (for given $z$) for larger $\theta_{R}$, and thus its amplitude is more attenuated. For example, if $c\rightarrow c_{\rm s}$, we have $\theta_{R}\rightarrow\pi/2$ and $\alpha_{1}=1/\tan\theta_{R}\rightarrow 0$ (from Snell’s law, we have $\sin(\theta_{R})=c/c_{\rm s}<1$). Since $K_{z}$ must increase for $\theta_{R}\rightarrow\pi/2$, we have $\tan\theta_{R}\rightarrow\infty$, and then one expects $K_{z}\propto 1/\alpha_{1}\rightarrow\infty$, and not $K_{z}\propto\alpha_{1}\rightarrow 0$ as in Shiokawa et al. (1990, 1989).